# CDEX Collaboration

### v1

@@ Abstract
It is worth noting that in Fig. 6, the CDEX-10 curve rises more slowly than the black dots (which is the current best $f_{PBH}$ limit obtained from cosmological observations Carr et al. (2010); Clark et al. (2018); Boudaud and Cirelli (2019); Coogan et al. (2021)). The XENON-1T results shown in Fig. 3 in Ref. Calabrese et al. (2022b) and in Fig. 4 in Ref. Li and Zhang (2022) exhibit the same trend. For example, the curve of ($m_{\chi}=$ 0.1 keV, $\sigma_{\chi e}=1\times 10^{-29}$ cm2) is out of the shaded part when $M_{PBH}\gtrsim 8\times 10^{16}$ g. This indicates that PBHeDM could be used to search for evaporating PBHs with large $M_{PBH}$s and imposes constraints on $f_{PBH}$s. DD experiments, such as XENON and CDEX, are expected to fill the gap in ($M_{PBH}$, $f_{PBH}$) for $M_{PBH}\sim\mathcal{O}(10^{18})$ g in Ref. Carr et al. (2021); Green and Kavanagh (2021) if the physical properties of $\chi$ can be clarified in the future.

### v2

@@ VII Discussion
All of the $f_{PBH}$s (including those from COMPTEL from other experiments) increase as $M_{PBH}$ increase because of the decreased Hawking temperature. For larger $M_{PBH}$, a lower energy threshold in the analysis is required. Due to the effect of Earth shielding, $f_{PBH}$ decreases with the decrease in $\sigma_{\chi e}$ between $\mathcal{O}(5\times 10^{-27})$ and $\mathcal{O}(1\times 10^{-28})$ cm2 in Fig. 6. Therefore, an accurate Earth shielding model rather than an approximation is required. This remains an open research area that will be the theme of our future efforts. Between $\mathcal{O}(1\times 10^{-28})$ and $\mathcal{O}(1\times 10^{-29})$ cm2, $f_{PBH}$ increases with the decrease in $\sigma_{\chi e}$.
It is worth noting that in Fig. 6, the CDEX-10 curve rises more slowly than the black dots (which is the current best $f_{PBH}$ limit obtained from cosmological observations Carr et al. (2010); Clark et al. (2018); Boudaud and Cirelli (2019); Coogan et al. (2021)). The XENON-1T’s results shown in Fig. 3 in Ref. Calabrese et al. (2022b) and in Fig. 4 in Ref. Li and Zhang (2022) exhibit the same trend. For example, the curve of ($m_{\chi}=$ 0.1 keV, $\sigma_{\chi e}=1\times 10^{-29}$ cm2) is out of the shaded part when $M_{PBH}\gtrsim 8\times 10^{16}$ g. This indicates that PBHeDM could be used to search for evaporating PBHs with large $M_{PBH}$s and imposes constraints on $f_{PBH}$s. DD experiments with low radiation background and low energy threshold, such as XENON and CDEX, are expected to fill the gap in ($M_{PBH}$, $f_{PBH}$) for $M_{PBH}\sim\mathcal{O}(10^{18})$ g in Ref. Carr et al. (2021); Green and Kavanagh (2021) if the physical properties of $\chi$ can be clarified in the future.
VIII Summary
PBHeDM is a novel non-thermal DM particle. In the presence of ($M_{PBH}$, $f_{PBH}$) compatible with present bounds, CDEX-10 experiment excludes $\sigma_{\chi e}\sim\mathcal{O}(10^{-28})$ cm2 for $\chi$ with a mass $m_{\chi}\lesssim$ 0.1 keV for $M_{PBH}$ ranging from $1\times 10^{15}$ g to $7\times 10^{16}$ g. This is a parameter space that previous works have not explored. In the future, if $\sigma_{\chi e}\left(m_{\chi}\right)$ is known, then DD experiments will have the potential to give limits on $f_{PBH}\left(M_{PBH}\right)$ for the larger $M_{PBH}$.
Acknowledgements